Find the equation of the straight line passing through (1, 2, 3) and perpendicular to the plane x + 2y – 5z + 9 = 0

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
equation of the line is
\frac{x-1}{1}=\frac{y-2}{2}=\frac{z-1}{-5}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty